Duracell Ltd claim that the life of their batteries in motorised soft toys is approximately normally distributed with a mean of 105.1 hours and a standard deviation of 18.2 hours. The best 77% of batteries would last beyond how many hours? Answer correct to 2 decimal places.
Answer #1

Answer = -1.54

Z = (x-μ) / σ

x = 77%

μ = 105.1%
σ = 18.2%

Z = (77 - 105.1) / 18.2 = -1.54

P(Z < -1.54) = 0.06179

-2 Area from a value (Use to compute p from Z) Value from an area (Use to compute Z for confidence intervals) Specify Paramet